OpenStudy (anonymous):

Let \[\log_{b}A=1 ; \log_{b}C=3 ; \log_{b}D=4 \] . What is the value of \[\log_{b}(\frac{ A^5C^2 }{ D^6 } )\] ? A. 0 B. 26 C. -13 D. There isn't enough information to answer the question.

OpenStudy (anonymous):

@Hero do you think you could please help me? :)

hero (hero):

Okay so basically what you do is use rules of logs to rewrite it as: \[\log_b(A^5) + \log_b(C^2) - \log(D^6)\] Then as: \[5\log_bA + 2\log_bC -6\log_bD\] Then substitute the appropriate values an finish solving.

OpenStudy (anonymous):

I keep getting logb (9/4096) but that doesn't seem right... what am I doing wrong?

hero (hero):

Then you end up with 5(1) + 2(3) - 6(4) = 5 + 6 - 24 = 11 - 24 = -13

OpenStudy (anonymous):

Oh I was substituting the wrong numbers in the wrong places! I see what you did now!

hero (hero):

I'm not an expert at these, but I think that should be correct.

OpenStudy (anonymous):

It was! thank you SO much :) do you know how to rewrite an expression of "log's" into a single logarithm ?

hero (hero):

I've done it before

OpenStudy (anonymous):

Rewrite the following expression as a single logarithm: 2log(x+3)+3log(x-7)-5log(x-2)-log(x^2) A. \[\log(\frac{ (x+3)^2(x-7)^3 }{ x^2(x-2)^5 })\] B. \[\log(\frac{ x^2(x+3)^2(x-7)^3 }{ (x-2)^5 })\] C. \[\log(\frac{ (x+3)^2(x-2)^5 }{ (x-7)^3x^2 })\] D. \[\log(\frac{ (x+3)^2(x-7)^3 }{ (x-2)^-5x^-2 })\]

hero (hero):

It's not that difficult at all

OpenStudy (anonymous):

what do i do? ive never done this

hero (hero):

First convert 2log(x+3)+3log(x-7)-5log(x-2)-log(x^2) to \[\log(x + 3)^2 + \log(x - 7)^3 - \log(x - 2)^5 - \log(x^2)\] Then convert that to \[\log((x + 3)^2(x - 7)^3) - \log\left(\frac{(x - 2)^5}{x^2}\right)\] Which then becomes: \[\log(a) - \log(b) = \log\left(\frac{a}{b}\right)\] Where a = \((x + 3)^2(x - 7)^3\) and b = \(\dfrac{(x - 2)^5}{x^2}\)

hero (hero):

So perform this division: \[a \div b = (x + 3)^2(x - 7)^3 \div \dfrac{(x - 2)^5}{x^2}\]

hero (hero):

Afterwards, put the log in front.

OpenStudy (anonymous):

I got A. Is that correct?

hero (hero):

Can you show your work please?

OpenStudy (anonymous):

well I thought that because you cant divide by an x you would just leave all of the functions however they were in the equation and since you cant fivide by x just add it to (x-2)^5 since it was a part of that function which would make it (x-2)^5+x^2

OpenStudy (anonymous):

right?

hero (hero):

So you didn't divide by hand. There's only one way to divide by a fraction. \[a \div \frac{b}{c} = a \times \frac{c}{b}\]

OpenStudy (anonymous):

I'm not in a place where I have writing utensils available so no, but I see what you did there, you divided using the reciprical/inverse right?

hero (hero):

What do you mean you're not in a place where you have writing utensils. You don't do math without such things.